Calculate the molarity of water if its density is 1000 kg/m3

Grade:upto college level

4 Answers

Navjyot Kalra
9 years ago
1 litre water = 1 kg i.e. 1000 g water (∵ d = 100 kg/m3)
≡ 1000/18 = 55.55 moles of water
So, molarity of water = 55.55M
Kriti Sharma
5 years ago
given that, 
             D= 1000kg/m^3
and         1m3= 1000 l 
so,       D= 1000kg/1000l = 1 kg/l
now,    1kg of water is present in 1l of solution 
now, no of moles of water=1000/18= 55.5 moles
so, molarity is 55.55 moles/l.
Lony Gohain
5 years ago
Density of water = 1000 kg/m^3
Molarity = no of moles of solute/volume of the                           solution
                =mass of solute/molar mass/volume of.                    Solution
Here solute = solution = water
 Molarity = mass of the solute / vol of soln / molar
                   Mass 
                = density of water/molar mass of water 
                = 1000/18   (molar mass of water= 18)
                = 55.56 mol/L
                   
 
Kushagra Madhukar
3 years ago
Dear student,
Please find the answer to your question.
 
1 litre water = 1 kg i.e. 1000 g water (∵ d = 100 kg/m3)
≡ 1000/18 = 55.55 moles of water
So, molarity of water = 55.55 M
 
Thanks and regards,
Kushagra